OpenStudy (anonymous):

simplify -2a(-7a^3+6)+(-2a)(6a^3-5)

OpenStudy (fibonaccichick666):

So can you tell me the order of operations?

OpenStudy (anonymous):

PEMDAS Parethesis Exponent Multiplications Division Addition Subtraction

OpenStudy (igreen):

Right..but in this case we only multiply.

OpenStudy (igreen):

And add.

OpenStudy (fibonaccichick666):

ok cool so, what is the first thing you should do?

OpenStudy (anonymous):

-2 times everything inside the first set of parentheses

OpenStudy (fibonaccichick666):

well you have -2a

OpenStudy (fibonaccichick666):

so though you could just do -2 first, why not get it all done in one shot?

OpenStudy (fibonaccichick666):

So can you try distributing the -2a for me?

OpenStudy (igreen):

Let's take the first part. \(-2a(-7a^3 + 6)\) Now we distribute -2a into the parenthesis. To do that, we just multiply it to every term inside the parenthesis. \(-2a \times -7a^3 =~?\) \(-2a \times 6 =~?\) Can you multiply those?

OpenStudy (anonymous):

-2a * -7a = 14a^4 -2a * 6 = -12a

OpenStudy (uri):

No kids I'm gonna explain

OpenStudy (uri):

It would be a^2 @TheYoungBlaze

OpenStudy (uri):

a*a=a^2

OpenStudy (anonymous):

I forgot to add the ^3 to it. It was actually -2a * -7a^3 as iGreen outlined.

OpenStudy (fibonaccichick666):

So @TheYoungBlaze , though you did the multiplication correctly, you wrote the steps incorrectly. So aanyways, can you do the second part of multiplication now?

OpenStudy (igreen):

YOu got it..

OpenStudy (uri):

My guys gotta go. My brother will help @iGreen :)

OpenStudy (igreen):

That gives us: \(-2a(-7a^3+6)+(-2a)(6a^3-5)\) \(14a^4 - 12a + (-2a)(6a^3 - 5)\) Now we distribute -2a into the parenthesis. Can you multiply these? \(-2a \times 6a^3\) \(-2a \times -5\)

OpenStudy (anonymous):

-2a * 6a^3 = -12a^4 -2a * -5 = -10a

OpenStudy (igreen):

Yep, that gives us: \(14a^4-12a - 12a^4 - 10a\) Can you combine like terms? \(14a^4 - 12a^4\) \(-12a - 10a\)

OpenStudy (anonymous):

14a^4 - 12a^4 = 2a^4 -12a - 10a = -22a

OpenStudy (igreen):

Yep, that gives us: \(2a^4 - 22a\) That's your final answer! :D

OpenStudy (anonymous):

So its 2a^4 - 22a?

OpenStudy (anonymous):

Hmm.. I double checked the original equation I posted and they're the same equation.. Idk where I went wrong?